The S&P 500 stock price closed at $98, $103, $107, $102, $111 over five successive weeks. What is the weekly standard deviation of the stock price? What is the weekly standard deviation of the returns assuming no dividend were paid?
Answer #1

1.
=sqrt((98-(98+103+107+102+111)/5)^2+(103-(98+103+107+102+111)/5)^2+(107-(98+103+107+102+111)/5)^2+(102-(98+103+107+102+111)/5)^2+(111-(98+103+107+102+111)/5)^2)/sqrt(5-1)=4.969909456

2.
=sqrt((111/102-1-(111/102-1+102/107-1+107/103-1+103/98-1))^2+(102/107-1-(111/102-1+102/107-1+107/103-1+103/98-1))^2+(107/103-1-(111/102-1+102/107-1+107/103-1+103/98-1))^2+(103/98-1-(111/102-1+102/107-1+107/103-1+103/98-1))^2)/sqrt(4-1)=12.727%